# Markovian Quantum Computation

### 04 Mar 2024 v1

@@ IV.3 Alternative approach
In the previous example reproducing Grover’s search algorithm, the transition term generates an all-to-all interaction with the identical transition strength. This makes the problem Hamiltonian completely featureless except for each energy level having a different degeneracy. While the absence of any local energy minimum is advantageous, the transition strength decreases superpolynomially with the number of qubits.
@@
The off-diagonal element ${}_{P}\langle z|H_{T}|z^{\prime}\rangle_{P}$ is then non-vanishing and equals one, independently of $N$, only when the Hamming distance between $z$ and $z^{\prime}$ is one. While the transition strength is maximum, local energy minima now become an issue. As the distance between two configurations is at most $N$, any state $|z\rangle_{P}$ can reach the ground state in at most $N$ transitions in principle. However, we do not know a general rule to take the right path for every state. For this reason, we take the direction of lowering the energy, even though it does not coincide with the direction of approaching the ground state. This mismatch is the origin of local energy minima.
@@
Note that every state $|z\rangle_{P}$ is linked to $N$ different states by transition. It is instructive to envisage this as an $N$-regular graph with $2^{N}$ vertices. Each vertex is endows with a potential determined by the problem Hamiltonian, and the transition occurs along the edges toward the direction of not increasing the potential. Once the population is trapped in a local potential minimum, a high-order transition is needed to get out of it.
@@
To elucidate this mechanism, consider an $n$-th order transition through the sequence $|z_{0}\rangle_{P}\rightarrow|z_{1}\rangle_{P}\rightarrow\cdots|z_{n}\rangle_{P}$. For the moment, suppose that $\alpha_{0}=1$ in Hamiltonian (2) and there is no cavity mode involved. For such a transition to occur, the first requirement is $E(z_{0})=E(z_{n})$. The perturbation theory yields the effective transition rate $\mathcal{O}[\lambda(\lambda/\Delta)^{n-1}]$. However, if the Stark shifts of $|z_{0}\rangle_{P}$ and $|z_{n}\rangle_{P}$ are different, the effective detuning breaks the condition $E(z_{0})=E(z_{n})$, suppressing the transition. Consequently, the most prominent transitions come from the cases wherein the energy differences satisfy the condition $E(z_{i+1})-E(z_{i})=E(z_{n-(i+1)})-E(z_{n-i})$ for all $i$. Now taking the cavity modes into account, the last condition can also be met when an intermediate transition $|z_{i}\rangle_{P}\rightarrow|z_{i+1}\rangle_{P}$ is replaced by $|z_{i}\rangle_{P}|0\rangle_{m}\rightarrow|z_{i+1}^{\prime}\rangle_{P}|1\rangle_{m}$ with $E(z_{i+1})=E(z_{i+1}^{\prime})+\omega_{m}$. This is always possible provided $E(z_{i+1}^{\prime})<E(z_{i+1})$ and the appropriate cavity mode exists. By incorporating multiple cavity modes, various tunnelling-out transition channels can appear. For instance, the lambda-type-like transitions characterized by $E(z_{1})-E(z_{0})=E(z_{n-1})-E(z_{n})+\omega_{m}$ with $n\geq 2$ are expected to play a crucial role in overcoming local minima.
@@
To illustrated the above idea, we consider a particular integer factoring algorithm (not to be confused with Shor’s algorithm [30, 22]). To be specific, we follow the integer multiplication procedure as in the elementary arithmetic and turn it into a combinatorial optimization problem. The problem Hamiltonian is constructed as follows. Consider a multiplication of two 3-bit integers $\bar{x}\equiv(x_{2}x_{1}x_{0})_{2}$ and $\bar{y}\equiv(y_{2}y_{1}y_{0})_{2}$ resulting in a 6-bit integer $\bar{z}\equiv(z_{5}z_{4}\cdots z_{0})_{2}$, incorporating additional four carry bits $\bar{c}\equiv(c_{3}c_{2}c_{1}c_{0})_{2}$. For a given value of $\bar{z}$, we turn each step of the calculation into an energy term in the problem Hamiltonian. Here, 10 qubits are needed to encode $\bar{x}$, $\bar{y}$, and $\bar{c}$, while $\bar{z}$ is hard-coded in the problem Hamiltonian. For example, the first energy term has value zero if $x_{0}y_{0}=z_{0}$ and $\Delta$ otherwise, the second has value zero if $x_{1}y_{0}+x_{0}y_{1}=(c_{0}z_{1})_{b}$ and $\Delta$ otherwise, and so on. The ground state, representing the solution of the factoring, has energy zero as it satisfies all the conditions, and has a two-fold degeneracy as $\bar{x}$ and $\bar{y}$ are interchangeable.
@@
Figure 2: Numerical simulation of MQC for integer factoring of 35. Dotted curves represent the energy with respect to the number of cooling cycles, averaged over $10^{3}$ samples, for (a) $\alpha_{0}=1$ and (b) $\alpha_{0}=0$. Gray curves represent typical individual trajectories.
@@
Fig. 2 shows the results of our numerical simulation for $\bar{z}=35$. We take the transition term as in Eq. (8) and set the parameters in Hamiltonian (2) as $M=3$, $\omega_{m}=m\Delta$, $\lambda=\Delta/10$, and $a_{0}=1$. The duration of a single cooling cycle is chosen to be $\pi/2\lambda$. The simulation is performed by exact diagonalization. Fig. 2(a) shows that the energy of the system vanishes in time, indicating that the system evolves into the ground state. As a comparison, we have changed $a_{0}$ to zero in Fig. 2(b) to test the effectiveness of the aforementioned mechanism for tunneling out from local minima. In the absence of such a mechanism, the system almost always evolves into local minima, which appear to be present in the first excited level.
@@
A remark is in order. Our choice of the parameter $\lambda=\Delta/10$, chosen to fulfill the requirement that the transition rate is much smaller than the energy difference, appears to be effective, as demonstrated in Fig. 2(a). However, it should be noted that the optimal choice of $\lambda$ depends on the characteristics and size of the problem. In general, smaller $\lambda$ is preferable as the collective effect enhance the transition rate. This effect is particularly important for combinatorial optimization problems, which are characterized by large degeneracies in the problem Hamiltonian. As the degeneracies of low energy levels are much smaller in general, the collective effect is also weaker in the low-energy part of the spectrum. This means that there exists a parametric regime of $\lambda$ where the heating is sufficiently avoided for the ground state, but not for higher energy states. On the other hand, by choosing smaller $\lambda$, leading to slower transitions, heating could be avoided throughout the entire energy spectrum. While the cooling works in both cases, it is unclear which one is more efficient.
@@ V Discussion
In this work, we have introduced the notion of Markovian quantum computation, along with a few illustrative examples. The idea is quite straightforward: cooling the system with a problem Hamiltonian through energy-lowering transitions, as in cavity cooling. An essential part here is to choose an appropriate transition term that generates off-diagonal elements in the diagonal basis of the problem Hamiltonian. There exists a trade-off between the connectivity and strength of the transitions. We have discussed two opposite limits when the ground state is the solution of a combinatorial optimization problem. In both cases, the running time generally increases superpolynomially with the system size. It is unclear if the performance can be substantially improved by interpolating between the two opposites.
MQC can also be used to find the ground state of general quantum Hamiltonians. The most important requirement is that the ground state is not heated during the process to make it a steady state. This is ensured when the ground state has a finite spectral gap to the excited state and this spectral gap can be resolved by a proper choice of the transition term and the parameters. Consequently, the inverse of the spectral gap determines the basic time scale of the computation. In principle, a finite number of cavity modes can cover the transitions across the entire spectrum of the problem Hamiltonian, where the required number of cavity modes is given roughly by the ratio of the energy range of the problem Hamiltonian to the characteristic transition rate. This means that the number of cavity modes increases polynomially with the system size for gapped Hamiltonians. However, choosing an appropriate transition term and estimating the resulting cooling time depend heavily on the system and are analytically challenging.

### 08 Mar 2024 v2

@@ IV.3 Alternative approach
In the previous example reproducing Grover’s search algorithm, the transition term generates an all-to-all interaction with the identical transition strength. This makes the problem Hamiltonian utterly featureless except for each energy level having a different degeneracy. While the absence of any local energy minimum is advantageous, the transition strength decreases superpolynomially with the number of qubits.
@@
The off-diagonal element ${}_{P}\langle z|H_{T}|z^{\prime}\rangle_{P}$ is then non-vanishing and equals one, independently of $N$, only when the Hamming distance between $z$ and $z^{\prime}$ is one. While the transition strength is maximum, local energy minima now become an issue. As the distance between two configurations is at most $N$, any state $|z\rangle_{P}$ can reach the ground state in at most $N$ transitions in principle. However, we do not know a general rule for every state to take the right path. For this reason, we take the direction of lowering the energy, even though it does not necessarily coincide with the direction of approaching the ground state. This mismatch is the origin of local energy minima.
@@
Note that every state $|z\rangle_{P}$ is linked to $N$ different states by transition. It is instructive to envisage this as an $N$-regular graph with $2^{N}$ vertices. Each vertex is endowed with a potential determined by the problem Hamiltonian, and the transition occurs along the edges toward the direction of not increasing the potential. Once the population is trapped in a local potential minimum, a high-order transition is needed to get out of it.
@@
To elucidate this mechanism, consider an $n$-th order transition through the sequence $|z_{0}\rangle_{P}\rightarrow|z_{1}\rangle_{P}\rightarrow\cdots|z_{n}\rangle_{P}$. For the moment, suppose that $\alpha_{0}=1$ in Hamiltonian (2) and there is no cavity mode involved. For such a transition to occur, the first requirement is $E(z_{0})=E(z_{n})$. The perturbation theory yields the effective transition rate $\mathcal{O}[\lambda(\lambda/\Delta)^{n-1}]$. However, if the Stark shifts of $|z_{0}\rangle_{P}$ and $|z_{n}\rangle_{P}$ are different, the effective detuning breaks the condition $E(z_{0})=E(z_{n})$, suppressing the transition. Consequently, the most prominent transitions come from the cases wherein the energy differences satisfy the condition $E(z_{i+1})-E(z_{i})=E(z_{n-(i+1)})-E(z_{n-i})$ for all $i$ (simply put, the transition diagram has a mirror symmetry). This rule is easily verifiable with simple numerical simulations, which we do not present here.
Now taking the cavity modes into account, the condition can also be met when an intermediate transition $|z_{i}\rangle_{P}\rightarrow|z_{i+1}\rangle_{P}$ is replaced by $|z_{i}\rangle_{P}|0\rangle_{m}\rightarrow|z_{i+1}^{\prime}\rangle_{P}|1\rangle_{m}$ with $E(z_{i+1})=E(z_{i+1}^{\prime})+\omega_{m}$. This is always possible provided $E(z_{i+1}^{\prime})<E(z_{i+1})$ and the appropriate cavity mode exists. By incorporating multiple cavity modes, various tunnelling-out transition channels can appear. For instance, the lambda-type-like transitions characterized by $E(z_{1})-E(z_{0})=E(z_{n-1})-E(z_{n})+\omega_{m}$ with $n\geq 2$ are expected to play an important role in overcoming local minima.
@@
To illustrate the above idea, we consider a particular integer factoring algorithm (not to be confused with Shor’s algorithm [38, 30]). Specifically, we follow the integer multiplication procedure as in the elementary arithmetic and turn it into a combinatorial optimization problem. The problem Hamiltonian is constructed as follows. Consider a multiplication of two 3-bit integers $\bar{x}\equiv(x_{2}x_{1}x_{0})_{2}$ and $\bar{y}\equiv(y_{2}y_{1}y_{0})_{2}$ resulting in a 6-bit integer $\bar{z}\equiv(z_{5}z_{4}\cdots z_{0})_{2}$, incorporating additional four carry bits $\bar{c}\equiv(c_{3}c_{2}c_{1}c_{0})_{2}$. For a given value of $\bar{z}$, we turn each calculation step into an energy term in the problem Hamiltonian. Here, 10 qubits are needed to encode $\bar{x}$, $\bar{y}$, and $\bar{c}$, while $\bar{z}$ is hard-coded in the problem Hamiltonian. For example, the first energy term has a value zero if $x_{0}y_{0}=z_{0}$ and $\Delta$ otherwise, the second has a value zero if $x_{1}y_{0}+x_{0}y_{1}=(c_{0}z_{1})_{2}$ and $\Delta$ otherwise, and so on. The ground state, representing the solution of the factoring, has energy zero as it satisfies all the conditions, and has a two-fold degeneracy as $\bar{x}$ and $\bar{y}$ are interchangeable.
@@
Figure 2: Numerical simulation of CQC for integer factoring of 35. Dotted curves represent the energy with respect to the number of cooling cycles, averaged over $10^{3}$ samples, for (a) $\alpha_{0}=1$ and (b) $\alpha_{0}=0$. Gray curves represent typical individual trajectories.
@@
Fig. 2 shows the results of our numerical simulation for $\bar{z}=35$. We take the transition term as in Eq. (8) and set the parameters in Hamiltonian (2) as $M=3$, $\omega_{m}=m\Delta$, $\lambda=\Delta/10$, and $a_{0}=1$. The duration of a single cooling cycle is chosen to be $\pi/2\lambda$. The simulation is performed by exact diagonalization. Fig. 2(a) shows that the energy of the system vanishes in time, indicating that the system evolves into the ground state. As a comparison, we have changed $a_{0}$ to zero for Fig. 2(b) to test the effectiveness of the mechanism for tunneling out from local minima. Without such a mechanism, the system almost always evolves into local minima, which appear to be present at the first excited level.
@@
A remark is in order. Our choice of the parameter $\lambda=\Delta/10$, chosen to fulfill the requirement that the transition rate is much smaller than the energy difference, appears effective, as demonstrated in Fig. 2(a). However, it should be noted that the optimal choice of $\lambda$ depends on the characteristics and size of the problem. In general, smaller $\lambda$ is preferable as the collective effect enhances the transition rate. This effect is particularly important for combinatorial optimization problems, which are characterized by large degeneracies in the problem Hamiltonian. As the degeneracies are generally very large in the middle of the energy spectrum, the collective effect is also stronger in that region. Consequently, there exists a parametric regime of $\lambda$ where the heating is sufficiently avoided for the ground state, but not for higher energy states. On the other hand, by choosing smaller $\lambda$, leading to slower transitions, one could reach the regime where heating is avoided throughout the entire energy spectrum. While the cooling is feasible in both cases, it is not clear which case is more efficient.
@@ V Discussion
In this work, we have introduced the notion of cooling-based quantum computation along with a few illustrative examples. An important part is to choose an appropriate transition term, which generates off-diagonal elements in the diagonal basis of the problem Hamiltonian. There exists a trade-off between the connectivity and strength of the transition. We have discussed two opposite limits when the ground state is the solution of a combinatorial optimization problem. In both cases, the running time generally increases superpolynomially with the system size. It is an open question whether the performance can be substantially improved by interpolating between the two opposites.
In fact, the idea of CQC can be used to prepare the ground states of general quantum Hamiltonians, although we focused on the analyzable cases. The most essential requirement is that the ground state should not be heated to make it a steady state. This is ensured when the ground state has a finite spectral gap to the excited state and this spectral gap is resolved. Consequently, the inverse of the spectral gap determines the basic time scale of the cooling, which is a natural consequence of the energy-time uncertainty principle. For gapped Hamiltonians, as the transition strength can be determined finitely without the risk of heating the ground state, a finite number of cavity modes can cover the transitions across the entire spectrum in principle, where the number is roughly given by the ratio of the whole energy range to the characteristic transition rate. This means that the number of cavity modes increases only polynomially with the system size for gapped Hamiltonians. However, choosing an appropriate transition term is heavily dependent on the Hamiltonian, hence analytically challenging.
An interesting open question is the precise relationship between CQC and AQC. In particular, the abundance and depth of local energy minima in CQC might be strongly correlated with the spectral gap during AQC [14, 32, 33]. If not, the two methods can differ in efficiency for particular ground-state findings. In fact, the transition terms $H_{T}$ in Eqs. (5) and (8) are the two conventional initial Hamiltonians in AQC. Consequently, when Trotter-decomposed, the infinitesimal unitary evolutions of CQC with Eqs. (5) and (8) are identical to those of AQC, although their arrangements and the inclusion or exclusion of cavities make differences. This might be a clue to understanding one of the two more deeply from inspecting the other.